OpenStudy (anonymous):

Find the distance between these points. C(0, 4), T(-6, -3)

Parth (parthkohli):

The distance formula: \(\Large \color{Black}{\Rightarrow \sqrt{(x_{2} - x_{1})^2 + (y_{2} - y_{1})^2} }\)

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow \sqrt{(-6)^2 + (-7)^2} }\) \(\Large \color{Black}{\Rightarrow \sqrt{36 + 49} }\) \(\Large \color{Black}{\Rightarrow \sqrt{85} }\)

OpenStudy (anonymous):

i love you parth thanks!!!!!!!!!!!!♥

OpenStudy (anonymous):

♥♥♥♥♥♥♥♥♥♥♥♥♥♥♥♥♥♥♥♥♥♥id give you more medals.

Parth (parthkohli):

Lol

Parth (parthkohli):

♥♥♥ :P

OpenStudy (anonymous):

:D